# The Voronoi Spherical CDF for Lattices and Linear Codes

### v2

@@ Abstract
For a lattice/linear code, we define the Voronoi spherical cumulative density function (CDF) as the CDF of the $\ell_{2}$-norm/Hamming weight of a random vector uniformly distributed over the Voronoi cell. Using the first moment method together with a simple application of Jensen’s inequality, we develop lower bounds on the expected Voronoi spherical CDF of a random lattice/linear code. Our bounds are valid for any finite dimension and are quite close to a trivial ball-based lower bound. They immediately translate to new non-asymptotic upper bounds on the normalized second moment and the error probability of a random lattice over the additive white Gaussian noise channel, as well as new non-asymptotic upper bounds on the Hamming distortion and the error probability of a random linear code over the binary symmetric channel. In particular, we show that for most lattices in $\mathbb{R}^{n}$ the second moment is greater than that of a Euclidean ball with the same covolume only by a $\left(1+O(\frac{1}{n})\right)$ multiplicative factor. Similarly, for most linear codes in $\mathbb{F}_{2}^{n}$ the expected Hamming distortion is greater than that of a corresponding Hamming ball only by an additive universal constant.
@@ I-D Acknowledgments
The author is grateful to Erik Agrell, Bruce Allen, Uri Erez, Bo’az Klartag, Yury Polyanskiy, Oded Regev and Barak Weiss for many valuable discussions that improved this paper. A special thanks is due to Erik Agrell for his valuable observation that our new upper bound in Theorem II.8 is close to the conjectured lower bound of Conway and Sloane, which led to the analysis in Subsection II-B1. We also thank the anonymous reviewers and associate editor for many helpful suggestions. This work was supported by ISF 1641/21 and ISF 2878/25.
@@ II-A Estimates for the expected Voronoi spherical CDF
$\underline{g}_{\mathrm{covering}}(r)=\begin{cases}1-e^{-(r/{r_{\mathrm{eff}}})^{n}}-23\cdot e^{-\eta/2}&(r/{r_{\mathrm{eff}}})^{n}<\eta/2\\ 1-24\cdot e^{-\eta/2}&\eta/2\leq(r/{r_{\mathrm{eff}}})^{n}<4n^{2}\eta\\ 1&(r/{r_{\mathrm{eff}}})^{n}\geq 4n^{2}\eta\end{cases}$
$\geq 1-\left(e^{-(r/{r_{\mathrm{eff}}})^{n}}+24e^{-\frac{\eta}{2}}\right)\mathds{1}\left\{\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}<4n^{2}\eta\right\}$
@@ II-B Bounds on the NSM
Finally, recalling that [68]*Section 3.241 for any $\nu>0$,
@@
$\leq\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}+23\cdot e^{-\eta/2}dr+\int_{r_{1}}^{r_{2}}24\cdot e^{-\eta/2}dr$ (45)
$\leq 24r_{2}e^{-\eta/2}+\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}dr.$ (46)
@@
$\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}dr$ $=\int_{0}^{r_{1}}e^{-cr^{\frac{n}{2}}}dr=\frac{2}{nc^{2/n}}\int_{0}^{cr_{1}^{n/2}}u^{\frac{2}{n}-1}e^{-u}du$ (47)
@@ II-B1 Gap to Conway and Sloane’s Conjectured Bound
Furthermore, in [71] an efficient algorithm with high numerical precision was developed for computing $f_{n}(x)$ in the range $x\in[n-1,n+1]$. Using this algorithm, Erik Agrell [72] has observed numerically that Zador’s upper bound (43) on the NSM is remarkably close to $\underline{G}^{\mathrm{CS}}_{n}$ for large $n$. See Figure 2 from [72]. Since our upper bound from Theorem II.8 is only $(1+c_{1}e^{-c_{2}n})$ greater than Zador’s upper bound, for universal $c_{1},c_{2}>0$, the same holds also for this bound. The next lemma, proved in Appendix C, validates the numerical findings of [72].
@@
Figure 2: Various bounds on the optimal NSM (due to [72].
@@ II-C Error probability of a random lattice
$P_{e}(L,\sigma^{2})=\Pr(\sigma Z\notin{\mathcal{V}_{L}})=1-\Pr(\sigma Z\in{\mathcal{V}_{L}})=1-\mu_{\sigma^{2}}({\mathcal{V}_{L}}).$ (57)
@@
$\mathbb{E}[P_{e}(L,\sigma^{2})]=P^{\mathrm{SP}}(\sigma^{2})+\left(\mathbb{E}[P_{e}(L,\sigma^{2})]-P^{\mathrm{SP}}(\sigma^{2})\right)$
$=P^{\mathrm{SP}}(\sigma^{2})+\left(\mu_{\sigma^{2}}({r_{\mathrm{eff}}}{\mathcal{B}})-{\mathbb{E}}[\mu_{\sigma^{2}}({\mathcal{V}_{L}})]\right)$
$=P^{\mathrm{SP}}(\sigma^{2})+{\mathbb{E}}_{\mu_{n}}[\Delta(L,\sigma^{2})]$
$=P^{\mathrm{SP}}(\sigma^{2})+{\mathbb{E}}_{W}\left[\frac{1}{X_{W}}\left(g_{\mathcal{B}}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)-{\mathbb{E}}_{\mu_{n}}\left[g_{L}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)\right]\right)\right]$
$=P^{\mathrm{SP}}(\sigma^{2})+{\mathbb{E}}_{W}\left[\frac{1}{X_{W}}\left(\min\{X_{W},1\}-{\mathbb{E}}_{\mu_{n}}\left[g_{L}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)\right]\right)\right]$ (65)
@@
Let $\tilde{Z}\sim\chi^{2}_{n}$ be a chi-squared random variable with $n$ degrees of freedom. The best known upper bound to date on ${\mathbb{E}}[P_{e}(L,\sigma^{2})]$,due to Poltyrev [7] and Ingber, Zamir, and Feder [8]*Theorem 2, see also [18]*Chapter 13.4, is the bound
@@
Figure 3: Computation of the sphere packing lower bound, the $P_{e}^{\mathrm{MLB}}(\sigma^{2})$ bound from [8]*Theorem 2, and the new upper bound from Theorem II.11. The bounds are plotted for $\sigma^{2}=\frac{0.95}{2\pi e}$ and $\sigma^{2}=\frac{0.98}{2\pi e}$.
@@ III-A Estimates for the expected Voronoi spherical CDF
$m_{{\mathcal{C}}}(\pi_{{\mathcal{C}}}({\mathcal{K}}))=\sum_{x\in{\mathcal{K}}}\frac{2^{k-n}}{|(x+{\mathcal{C}})\cap{\mathcal{K}}|}=\sum_{x\in{\mathcal{K}}}\frac{2^{k-n}}{1+|({\mathcal{C}}\setminus\{0\})\cap({\mathcal{K}}-x)|}.$ (82)
@@
${\mathbb{E}}[Q_{\mathcal{C}}(r)]={\mathbb{E}}[\Pr(|U_{\mathcal{C}}|\leq r)]\geq\frac{2^{k-n}|{\mathcal{B}}_{n,r}|}{1+2^{k-n}|{\mathcal{B}}_{n,r}|}=\frac{2^{k-n}V_{n,r}}{1+2^{k-n}V_{n,r}},$ (90)
@@ III-B Hamming distortion of a linear code
$=\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}\frac{x_{r}^{2}}{1+x_{r}}+\sum_{r={r_{\mathrm{eff}}}(n,k)}^{n-1}\frac{1}{1+x_{r}},$ (100)
@@ III-B1 Lossy compression of binary symmetric source under Hamming distortion
$n{\mathbb{E}}[D]$ $=\sum_{r=0}^{n}\left(1-2^{-n}V_{n,r}\right)^{2^{k}}\leq\sum_{r=0}^{n}\left(1-2^{k-n}V_{n,r}\right)$
@@ III-C Error probability of a linear code
$\mathbb{E}[P_{e}({\mathcal{C}},p)]$ $=P^{\mathrm{SP}}_{n,k}(p)+\left(\mathbb{E}[P_{e}({\mathcal{C}},p)]-P^{\mathrm{SP}}_{n,k}(p)\right)$
@@
${\mathbb{E}}[\mu_{p}(\mathcal{V}_{\mathcal{C}})]\geq 2^{n-k}\left(\frac{p}{1-p}\cdot p^{n}+\frac{1-2p}{1-p}{\mathbb{E}}\left[\frac{2^{k-n}V_{n,|Z|}}{{{n}\choose{|Z|}}\left(1+2^{k-n}V_{n,|Z|}\right)}\right]\right).$ (116)
@@
$\Delta^{p}_{n,k}$ ${\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\mu_{p}(\bar{{\mathcal{B}}}^{n,k})-{\mathbb{E}}[\mu_{p}(\mathcal{V}_{\mathcal{C}})]$
$\leq\frac{1-2p}{1-p}2^{n-k}{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\min\{x_{|Z|},1\}-\frac{x_{|Z|}}{1+x_{|Z|}}\right)\right]$
$=\frac{1-2p}{1-p}2^{n-k}{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\frac{\min\{x^{2}_{|Z|},1\}}{1+x_{|Z|}}\right)\right]$
$=\frac{1-2p}{1-p}2^{n-k}{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\frac{\min\{2^{k-n}V_{n,|Z|}\cdot x_{|Z|},2^{k-n}V_{n,|Z|}\cdot x^{-1}_{|Z|}\}}{1+x_{|Z|}}\right)\right]$
@@
Figure 4: Computation of the sphere packing lower bound, the RCU bound from [10]*Theorem 33, and the new upper bound from Theorem III.11. The bounds are plotted for $p=0.07$ and $p=0.1$, even values of $n$ and $k=n/2$.
@@
The benchmark upper bound to date for the binary symmetric channel is the so-called random coding upper bound (RCU bound) from [9] and [10]*Theorem 33, which implies that for a random linear code
@@ Appendix A Proofs of claims for lattices
${r_{\mathrm{pack}}}(L)=\sup\{r\penalty 10000\ :\penalty 10000\ r{\mathcal{B}}\subset{\mathcal{V}_{L}}\}=\sup\left\{r\penalty 10000\ :\penalty 10000\ g_{L}(r)=\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}\right\},$ (119)
@@
${\mathbb{E}}\|U_{L}\|^{2}=\int_{r=0}^{\infty}\Pr(\|U_{L}\|^{2}>r)dr=\int_{r=0}^{\infty}1-\Pr(\|U_{L}\|\leq\sqrt{r})dr.$ (121)
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \cdot\int_{0}^{\infty}\frac{w^{\frac{n+2}{2}-1}e^{-\frac{w}{2}}}{2^{\frac{n+2}{2}}\Gamma\left(\frac{n+2}{2}\right)}\frac{\Pr\left(\|U\|_{2}\leq\sqrt{\sigma^{2}w}\right)}{w^{\frac{n}{2}}}dw$ (127)
@@
$\varepsilon(J,L)=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\mathds{1}\{|L\cap(x+J)|=\emptyset\}dx.$ (131)
@@
${\mathbb{E}}_{\mu_{n}}[\varepsilon(J,L)]=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}{\mathbb{E}}_{\mu_{n}}[\mathds{1}\{|L\cap(x+J)|=\emptyset\}]dx$ (132)
$=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\mu_{n}\left\{L\in{\mathcal{L}}_{n}\penalty 10000\ :\penalty 10000\ L\penalty 10000\ \text{is}\penalty 10000\ (x+J)-\text{admissible}\right\}dx$ (133)
$=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}e^{-V}(1-R)dx$ (134)
@@
${\mathbb{E}}_{\mu_{n}}[g_{L}(r)]$ $=\Pr(L\in\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\in\mathcal{E}_{\eta}]+\Pr(L\notin\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\notin\mathcal{E}_{\eta}]$
$=\Pr(L\in\mathcal{E}_{\eta}){\mathbb{E}}_{\tilde{\mu}_{n}}[g_{L}(r)]+\Pr(L\notin\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\notin\mathcal{E}_{\eta}]$
@@
We use the standard asymptotic expansions [77, 6.141, 6.133] [78, Chapter 6.3]
$\log\Gamma(z)$ $=\Bigl(z-\tfrac{1}{2}\Bigr)\log z-z+\tfrac{1}{2}\log(2\pi)+O(z^{-1}),\qquad z\to\infty,$ (144)
$\log\Gamma(1+\varepsilon)$ $=-\gamma\,\varepsilon+O(\varepsilon^{2}),\qquad\varepsilon\to 0,$ (145)
$H_{n+2}$ $=\log n+\gamma+O(n^{-1}),\qquad n\to\infty.$ (146)
@@
$\log\underline{G}^{\mathrm{CS}}_{n}$ $=\log\!\Bigl(\frac{n+3-2H_{n+2}}{4n(n+1)}\Bigr)+\frac{\log(n+1)}{n}+\frac{4}{n}\log(n!)+\frac{2}{n}\log f_{n}(n)$
@@
$\frac{n+3-2H_{n+2}}{n+1}=\frac{n+1-2(H_{n+2}-1)}{n+1}=1-\frac{2(H_{n+2}-1)}{n+1}$
$=1-2\frac{\log(n)+(\gamma-1)+O(n^{-1})}{n+1}$
@@
$=-\log(4n)+\log\left(1-\frac{2\log{n}+2(\gamma-1)}{n}+O\left(\frac{\log n}{n^{2}}\right)\right)$
$=-\log(4n)-\frac{2\log{n}+2(\gamma-1)}{n}-\frac{2\log^{2}(n)}{n}+O\left(\frac{\log(n)}{n^{2}}\right)$ (153)
@@
$\log\underline{G}^{\mathrm{CS}}_{n}=-\log(2\pi)-1+\frac{\log(\pi n)-2\gamma}{n}-\frac{2\log^{2}n}{n^{2}}+O\left(\frac{\log n}{n^{2}}\right).$ (155)
Combining (150) with (155) gives

### v3

@@ Abstract
For a lattice/linear code, we define the Voronoi spherical cumulative density function (CDF) as the CDF of the $\ell_{2}$-norm/Hamming weight of a random vector uniformly distributed over the Voronoi cell. Using the first moment method together with a simple application of Jensen’s inequality, we develop lower bounds on the expected Voronoi spherical CDF of a random lattice/linear code. Our bounds are valid for any finite dimension and are quite close to a ball-based lower bound. They immediately translate to new non-asymptotic upper bounds on the normalized second moment and the error probability of a random lattice over the additive white Gaussian noise channel, as well as new non-asymptotic upper bounds on the Hamming distortion and the error probability of a random linear code over the binary symmetric channel. In particular, we show that for most lattices in $\mathbb{R}^{n}$ the second moment is greater than that of a Euclidean ball with the same covolume only by a $\left(1+O(\frac{1}{n})\right)$ multiplicative factor. Similarly, for most linear codes in $\mathbb{F}_{2}^{n}$ the expected Hamming distortion is greater than that of a corresponding Hamming ball only by an additive universal constant.
@@ I-D Acknowledgments
The author is grateful to Erik Agrell, Bruce Allen, Uri Erez, Bo’az Klartag, Yury Polyanskiy, Oded Regev and Barak Weiss for many valuable discussions that improved this paper. A special thanks is due to Erik Agrell for his valuable observation that our new upper bound in Theorem II.8 is close to the conjectured lower bound of Conway and Sloane, which led to the analysis in Subsection II-B1. We also thank the anonymous reviewers and associate editor for many helpful suggestions.
@@ II-A Estimates for the expected Voronoi spherical CDF
$\underline{g}_{\mathrm{covering}}(r)$
$=\begin{cases}1-e^{-\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}}-23\cdot e^{-\frac{\eta}{2}}&\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}<\frac{\eta}{2}\\ 1-24\cdot e^{-\frac{\eta}{2}}&\frac{\eta}{2}\leq\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}<4n^{2}\eta\\ 1&\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}\geq 4n^{2}\eta\end{cases}$
$\geq 1-\left(e^{-\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}}+24e^{-\frac{\eta}{2}}\right)\mathds{1}\left\{\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}<4n^{2}\eta\right\}$
@@ II-B Bounds on the NSM
Finally, recalling that [68, Section 3.241] for any $\nu>0$,
@@
$\leq\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}+23\cdot e^{-\eta/2}dr+\int_{r_{1}}^{r_{2}}24\cdot e^{-\frac{\eta}{2}}dr$ (45)
$\leq 24r_{2}e^{-\frac{\eta}{2}}+\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}dr.$ (46)
@@
$\int_{0}^{r_{1}}e^{-\left(\frac{r}{{r_{\mathrm{eff}}}^{2}}\right)^{\frac{n}{2}}}dr$ $=\int_{0}^{r_{1}}e^{-cr^{\frac{n}{2}}}dr$ (47)
$=\frac{2}{nc^{2/n}}\int_{0}^{cr_{1}^{n/2}}u^{\frac{2}{n}-1}e^{-u}du$ (48)
@@ II-B1 Gap to Conway and Sloane’s Conjectured Bound
Furthermore, in [71] an efficient algorithm with high numerical precision was developed for computing $f_{n}(x)$ in the range $x\in[n-1,n+1]$. Using this algorithm, Erik Agrell [72] has observed numerically that Zador’s upper bound (43) on the NSM is remarkably close to $\underline{G}^{\mathrm{CS}}_{n}$ for large $n$. See Figure 2,reproduced from [72] (see also [63, Figure 1] for a similar figure). Since our upper bound from Theorem II.8 is only $(1+c_{1}e^{-c_{2}n})$ greater than Zador’s upper bound, for universal $c_{1},c_{2}>0$, the same holds also for this bound. The next lemma, proved in Appendix C, validates the numerical findings of [72].
@@
Figure 2: Various bounds on the optimal NSM (due to [72]).
@@ II-C Error probability of a random lattice
$P_{e}(L,\sigma^{2})$ $=\Pr(\sigma Z\notin{\mathcal{V}_{L}})=1-\Pr(\sigma Z\in{\mathcal{V}_{L}})$ (58)
$=1-\mu_{\sigma^{2}}({\mathcal{V}_{L}}).$ (59)
@@
$\mathbb{E}[P_{e}(L,\sigma^{2})]=P^{\mathrm{SP}}(\sigma^{2})+\left(\mathbb{E}[P_{e}(L,\sigma^{2})]-P^{\mathrm{SP}}(\sigma^{2})\right)$ (67)
$=P^{\mathrm{SP}}(\sigma^{2})+\left(\mu_{\sigma^{2}}({r_{\mathrm{eff}}}{\mathcal{B}})-{\mathbb{E}}[\mu_{\sigma^{2}}({\mathcal{V}_{L}})]\right)$ (68)
$=P^{\mathrm{SP}}(\sigma^{2})+{\mathbb{E}}_{\mu_{n}}[\Delta(L,\sigma^{2})]$ (69)
$=P^{\mathrm{SP}}(\sigma^{2})$
$+{\mathbb{E}}_{W}\left[\frac{1}{X_{W}}\left(g_{\mathcal{B}}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)-{\mathbb{E}}_{\mu_{n}}\left[g_{L}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)\right]\right)\right]$ (70)
$=P^{\mathrm{SP}}(\sigma^{2})$
$+{\mathbb{E}}_{W}\left[\frac{1}{X_{W}}\left(\min\{X_{W},1\}-{\mathbb{E}}_{\mu_{n}}\left[g_{L}\left({r_{\mathrm{eff}}}X_{W}^{1/n}\right)\right]\right)\right]$ (71)
@@
Let $\tilde{Z}\sim\chi^{2}_{n}$ be a chi-squared random variable with $n$ degrees of freedom. The best known upper bound to date on ${\mathbb{E}}[P_{e}(L,\sigma^{2})]$,due to Poltyrev [7] and Ingber, Zamir, and Feder [8, Theorem 2], see also [18, Chapter 13.4], is the bound
@@
Figure 3: Computation of the sphere packing lower bound, the $P_{e}^{\mathrm{MLB}}(\sigma^{2})$ bound from [8, Theorem 2], and the new upper bound from Theorem II.11. The bounds are plotted for $\sigma^{2}=\frac{0.95}{2\pi e}$ and $\sigma^{2}=\frac{0.98}{2\pi e}$.
@@ III-A Estimates for the expected Voronoi spherical CDF
$m_{{\mathcal{C}}}(\pi_{{\mathcal{C}}}({\mathcal{K}}))$ $=\sum_{x\in{\mathcal{K}}}\frac{2^{k-n}}{|(x+{\mathcal{C}})\cap{\mathcal{K}}|}$ (88)
$=\sum_{x\in{\mathcal{K}}}\frac{2^{k-n}}{1+|({\mathcal{C}}\setminus\{0\})\cap({\mathcal{K}}-x)|}.$ (89)
@@
${\mathbb{E}}[Q_{\mathcal{C}}(r)]={\mathbb{E}}[\Pr(|U_{\mathcal{C}}|\leq r)]$ $\geq\frac{2^{k-n}|{\mathcal{B}}_{n,r}|}{1+2^{k-n}|{\mathcal{B}}_{n,r}|}$ (97)
$=\frac{2^{k-n}V_{n,r}}{1+2^{k-n}V_{n,r}},$ (98)
@@ III-B Hamming distortion of a linear code
$\leq\sum_{r=0}^{{r_{\mathrm{eff}}}(n,k)-1}\frac{x_{r}^{2}}{1+x_{r}}+\sum_{r={r_{\mathrm{eff}}}(n,k)}^{n-1}\frac{1}{1+x_{r}},$ (108)
@@ III-B1 Lossy compression of binary symmetric source under Hamming distortion
$n{\mathbb{E}}[D]$ $=\sum_{r=0}^{n}\left(1-2^{-n}V_{n,r}\right)^{2^{k}}\leq\sum_{r=0}^{n}\left(1-2^{k-n}V_{n,r}\right)$ (116)
@@ III-C Error probability of a linear code
$\mathbb{E}[P_{e}({\mathcal{C}},p)]$ $=P^{\mathrm{SP}}_{n,k}(p)+\left(\mathbb{E}[P_{e}({\mathcal{C}},p)]-P^{\mathrm{SP}}_{n,k}(p)\right)$ (123)
@@
${\mathbb{E}}[\mu_{p}(\mathcal{V}_{\mathcal{C}})]$
$\geq 2^{n-k}\left(\frac{p}{1-p}p^{n}+\frac{1-2p}{1-p}{\mathbb{E}}\left[\frac{2^{k-n}V_{n,|Z|}}{{{n}\choose{|Z|}}\left(1+2^{k-n}V_{n,|Z|}\right)}\right]\right).$ (126)
@@
$\Delta^{p}_{n,k}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\mu_{p}(\bar{{\mathcal{B}}}^{n,k})-{\mathbb{E}}[\mu_{p}(\mathcal{V}_{\mathcal{C}})]$ (127)
$\leq\frac{1-2p}{1-p}2^{n-k}{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\min\{x_{|Z|},1\}-\frac{x_{|Z|}}{1+x_{|Z|}}\right)\right]$ (128)
$=\frac{1-2p}{1-p}2^{n-k}{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\frac{\min\{x^{2}_{|Z|},1\}}{1+x_{|Z|}}\right)\right]$ (129)
$=\frac{1-2p}{1-p}2^{n-k}$
$\cdot{\mathbb{E}}\left[\frac{1}{{{n}\choose{|Z|}}}\left(\frac{\min\{2^{k-n}V_{n,|Z|}\cdot x_{|Z|},2^{k-n}V_{n,|Z|}\cdot x^{-1}_{|Z|}\}}{1+x_{|Z|}}\right)\right]$ (130)
@@
Figure 4: Computation of the sphere packing lower bound, the RCU bound from [10, Theorem 33], and the new upper bound from Theorem III.11. The bounds are plotted for $p=0.07$ and $p=0.1$, even values of $n$ and $k=n/2$.
@@
The benchmark upper bound to date for the binary symmetric channel is the so-called random coding upper bound (RCU bound) from [9] and [10, Theorem 33], which implies that for a random linear code
@@ Appendix A Proofs of claims for lattices
${r_{\mathrm{pack}}}(L)$ $=\sup\{r\penalty 10000\ :\penalty 10000\ r{\mathcal{B}}\subset{\mathcal{V}_{L}}\}$ (133)
$=\sup\left\{r\penalty 10000\ :\penalty 10000\ g_{L}(r)=\left(\frac{r}{{r_{\mathrm{eff}}}}\right)^{n}\right\},$ (134)
@@
${\mathbb{E}}\|U_{L}\|^{2}$ $=\int_{r=0}^{\infty}\Pr(\|U_{L}\|^{2}>r)dr$ (136)
$=\int_{r=0}^{\infty}1-\Pr(\|U_{L}\|\leq\sqrt{r})dr.$ (137)
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \cdot\int_{0}^{\infty}\frac{w^{\frac{n+2}{2}-1}e^{-\frac{w}{2}}}{2^{\frac{n+2}{2}}\Gamma\left(\frac{n+2}{2}\right)}\frac{\Pr\left(\|U\|_{2}\leq\sqrt{\sigma^{2}w}\right)}{w^{\frac{n}{2}}}dw$ (143)
@@
$\varepsilon(J,L)=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\hskip-28.45274pt\mathds{1}\{|L\cap(x+J)|=\emptyset\}dx.$ (147)
@@
${\mathbb{E}}_{\mu_{n}}[\varepsilon(J,L)]$ (148)
$=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\hskip-28.45274pt{\mathbb{E}}_{\mu_{n}}[\mathds{1}\{|L\cap(x+J)|=\emptyset\}]dx$ (149)
$=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\hskip-34.1433pt\mu_{n}\left\{L\in{\mathcal{L}}_{n}\penalty 10000\ :\penalty 10000\ L\penalty 10000\ \text{is}\penalty 10000\ (x+J)-\text{admissible}\right\}dx$ (150)
$=\lim_{a\to\infty}\frac{1}{a^{n}}\int_{x\in\mathrm{CUBE}(a)}\hskip-28.45274pte^{-V}(1-R)dx$ (151)
@@
${\mathbb{E}}_{\mu_{n}}[g_{L}(r)]$
$=\Pr(L\in\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\in\mathcal{E}_{\eta}]+\Pr(L\notin\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\notin\mathcal{E}_{\eta}]$ (157)
$=\Pr(L\in\mathcal{E}_{\eta}){\mathbb{E}}_{\tilde{\mu}_{n}}[g_{L}(r)]+\Pr(L\notin\mathcal{E}_{\eta}){\mathbb{E}}[g_{L}(r)|L\notin\mathcal{E}_{\eta}]$ (158)
@@
We use the standard asymptotic expansions for large $z$, small $\varepsilon$ and large $n$ [77, 6.141, 6.133] [78, Chapter 6.3]
$\log\Gamma(z)$ $=\Bigl(z-\tfrac{1}{2}\Bigr)\log z-z+\tfrac{1}{2}\log(2\pi)+O(z^{-1}),$ (163)
$\log\Gamma(1+\varepsilon)$ $=-\gamma\,\varepsilon+O(\varepsilon^{2}),$ (164)
$H_{n+2}$ $=\log n+\gamma+O(n^{-1}).$ (165)
@@
$\log\underline{G}^{\mathrm{CS}}_{n}=\log\!\Bigl(\frac{n+3-2H_{n+2}}{4n(n+1)}\Bigr)+\frac{\log(n+1)}{n}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{4}{n}\log(n!)+\frac{2}{n}\log f_{n}(n)$ (170)
@@
$\frac{n+3-2H_{n+2}}{n+1}=\frac{n+1-2(H_{n+2}-1)}{n+1}$ (172)
$=1-\frac{2(H_{n+2}-1)}{n+1}$ (173)
$=1-2\frac{\log(n)+(\gamma-1)+O(n^{-1})}{n+1}$ (174)
@@
$=-\log(4n)$
$+\log\left(1-\frac{2\log{n}+2(\gamma-1)}{n}+O\left(\frac{\log n}{n^{2}}\right)\right)$ (176)
$=-\log(4n)$
$-\frac{2\log{n}+2(\gamma-1)}{n}-\frac{2\log^{2}(n)}{n}+O\left(\frac{\log(n)}{n^{2}}\right)$ (177)
@@
$\log\underline{G}^{\mathrm{CS}}_{n}=-\log(2\pi)-1+\frac{\log(\pi n)-2\gamma}{n}$
$-\frac{2\log^{2}n}{n^{2}}+O\left(\frac{\log n}{n^{2}}\right).$ (179)
Combining (169) with (C) gives